Solve the differential equation (D^2+4)y=cosec2x in the method of variation of parameters
1
Expert's answer
2020-08-03T18:49:20-0400

General solution of "y''+4y=0" is "y=A\cos 2x+B\sin 2x".

Let "y=A(x)\cos 2x+B(x)\sin 2x" and "A'(x)\cos 2x+B'(x)\sin2x=0". Then "y'=-2A(x)\sin 2x+2B(x)\cos 2x" and so "y''=-2A'(x)\sin 2x+2B'(x)\cos 2x-4A(x)\cos 2x-4B(x)\sin 2x".

We have "\frac{1}{\sin 2x}=y''+4y=-2A'(x)\sin 2x+2B'(x)\cos 2x". Also we have "A'(x)\cos 2x+B'(x)\sin2x=0".


"-1=0\cdot 2\cos 2x-\frac{1}{\sin 2x}\sin 2x="

"=2(A'(x)\cos 2x+B'(x)\sin2x)\cos 2x-"

"-(-2A'(x)\sin 2x+2B'(x)\cos 2x)\sin 2x=2A'(x)", so "A'(x)=-\frac{1}{2}" and "A(x)=-\frac{1}{2}x+A"


Since "0=A'(x)\cos 2x+B'(x)\sin2x=-\frac{1}{2}\cos 2x+B'(x)\sin2x", we have "B'(x)=\frac{1}{2}\frac{\cos 2x}{\sin 2x}". Then "dB=\frac{1}{2}\frac{\cos 2x}{\sin 2x}=\frac{1}{4}\frac{d\sin 2x}{\sin 2x}=\frac{1}{4}d\ln|\sin 2x|" and "B(x)=\frac{1}{4}\ln|\sin 2x|+B"


We obtain "y=\left(-\frac{1}{2}x+A\right)\cos 2x+\left(\frac{1}{4}\ln|\sin 2x|+B\right)\sin 2x"

Answer: "y=\left(-\frac{1}{2}x+A\right)\cos 2x+\left(\frac{1}{4}\ln|\sin 2x|+B\right)\sin 2x"